OpenStudy (cherie_magee):

Solve the following system of equations by the substitution method. 10x + 10y = 1 x = y - 3 What is the value of y? -29/20 31/20 -3/2

OpenStudy (fifciol):

substitute y in place of x in first equation

OpenStudy (fifciol):

10(y-3) +10y=1 solve for y

OpenStudy (cherie_magee):

i need a lil more help then that @Fifciol

OpenStudy (ybarrap):

@Fifciol is right. X is already solved for you, it's just x=y-3. Next you need to find y. That's what he did.

OpenStudy (anonymous):

do you need help solving \[10(y-3) +10y=1\]?

OpenStudy (anonymous):

10(y-3)+10y=1 =>10y-30+10y=1 =>20y=1+30 =>y=31/20